OpenStudy (anonymous):

If 2W=3X and W+Z=X+Y then what is the value of Y?

OpenStudy (helder_edwin):

u have (1) W+Z=X+Y and (2) 2W=3X. if u multiply (1) by 2 u get 2W+2Z=2X+2Y and replacing (2) into this u get 3X+2Z=2X+2Y X+2Z=2Y Y=1/2(X+2Z)=1/2X+Z

OpenStudy (anonymous):

I forgot to mention that Y must be in terms of W...

OpenStudy (helder_edwin):

so multiply (1) by 3 and get 3W+3Z=3X+3Y and replacing (2) into this 3W+3Z=2W+3Y W+3Z=3Y Y=1/3(W+3Z)=1/3W+Z

OpenStudy (anonymous):